Suppose that the labor market for truck drivers is defined as such:

Labor Demand:Wage=21-0.01Q

Labor Supply:Wage=0.01Q

Calculate the following:

Market equilibrium wage for truck drivers Market equilibrium quantity of workers for truck drivers.

Consumer Surplus (what does this mean in the context of the Labor market?)

Producer Surplus (what does this mean in the context of the Labor market?)
Answer #1

Setting QD=QS

21-0.01Q=0.01Q

21=0.01Q+0.01Q

Q=21/0.02 = 1050

P=21-0.01(1050) = 10.5

CS = 0.5*1050*(21-10.5) = 5512.5

Consumer surplus is the area which shows the welfare of the employers in the labor market

PS = 0.5*1050*(10.5-0) = 5512.5

Producer surplus is the area which shows the welfare of the workers in the labor market.